# Disentangled Representation Learning with the Gromov-Monge Gap

### v2

@@ B.2 On rescaling the costs matrices in the entropic GW solver
$\begin{split}&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-2\langle\frac{\mathbf{C}_{\mathcal{X}}}{\texttt{stat}(\mathbf{C}_{\mathcal{X}})}\mathbf{P}\frac{\mathbf{C}_{\mathcal{Y}}}{\texttt{stat}(\mathbf{C}_{\mathcal{Y}})},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,\\ \Leftrightarrow\quad&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-\frac{2}{{\texttt{stat}(\mathbf{C}_{\mathcal{X}})}\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})}\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\varepsilon H(\mathbf{P})\,\\ \Leftrightarrow\quad&\tilde{\mathbf{P}}^{\star}\in\operatorname*{arg\,min}_{\mathbf{P}\in U_{n}}\,-2\langle\mathbf{C}_{\mathcal{X}}\mathbf{P}\mathbf{C}_{\mathcal{Y}},\mathbf{P}\rangle-\texttt{stat}(\mathbf{C}_{\mathcal{X}})\cdot\texttt{stat}(\mathbf{C}_{\mathcal{Y}})\cdot\varepsilon H(\mathbf{P})\,,\end{split}$ (13)
@@ B.3 Positivity of the Entropic GMG estimator
$\begin{split}\mathrm{GMG}_{{r}_{n},\varepsilon}(T)&=\mathrm{DST}_{r_{n}}(T)-\mathrm{GW}_{\varepsilon}({r}_{n},T\sharp{r}_{n})\,\\ &=\mathrm{DST}_{r_{n}}(T)-\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})\,,\end{split}$
For any coupling $\mathbf{P}\in U_{n}$, since $-\varepsilon H(\mathbf{P})=-\varepsilon\sum_{i,j=1}^{n}\mathbf{P}_{ij}\log(\mathbf{P}_{ij})<0$, one has:
@@
B.5 Proofs of Thm. 3.5
See 3.5
Before proving Thm. 3.5, we first demonstrate some technical results that will be useful later.
@@
Let ${p}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ and ${q}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{y}_{i}}$ two uniform, empirical measures, supported on the same number of points. We denote by $P_{n}=\{\mathbf{P}\in\mathbb{R}^{n\times n},\exists\sigma\in\mathcal{S}_{n},\mathbf{P}_{ij}:=\delta_{j,\sigma(i)}\}$ the set set of permutation matrices. Assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels (see A.1). Then, the GM and GW formulations coincide, in the sense that we can restrict the GW problem to permutations, namely
@@
As we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels, from (Séjourné et al., 2023, Theorem. 3), $\mathbf{P}^{\star}$ also solves:
@@
$\begin{split}\mathrm{GW}^{\langle\cdot,\cdot\rangle}(p,q)=&\int_{\mathcal{X}\times\mathcal{X}}\langle\mathbf{x},\mathbf{x}^{\prime}\rangle\mathop{}\!\mathrm{d}p(\mathbf{x})\mathop{}\!\mathrm{d}p(\mathbf{x}^{\prime})+\int_{\mathcal{Y}\times\mathcal{Y}}\langle\mathbf{y},\mathbf{y}^{\prime}\rangle\mathop{}\!\mathrm{d}q(\mathbf{y})\mathop{}\!\mathrm{d}q(\mathbf{y}^{\prime})\\ +&\min_{\mathbf{M}\in\mathcal{M}}\min_{\pi\in\Pi(p,q)}\int_{\mathcal{X}\times\mathcal{Y}}-4\langle\mathbf{M}\mathbf{x},\mathbf{y}\rangle\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})+4\|\mathbf{M}\|_{2}^{2}\,,\end{split}$ (19)
@@
We use the fact that a twice continuously differentiable function is convex i.f.f. its hessian is positive semi-definite (Boyd and Vandenberghe, 2004, §(3.1.4)). Therefore, $f_{\mathbf{A}}$ is convex i.f.f. $\nabla^{2}f_{\mathbf{A}}=\mathbf{A}\geq 0$. If $\lambda_{\min}(\mathbf{A})\geq 0$, then $\mathbf{A}\geq 0$ so $f_{\mathbf{A}}$ is convex, i.e. 00-weakly convex. Otherwise, $f_{\mathbf{A}}-\tfrac{1}{2}\lambda_{\min}(\mathbf{A})\|\cdot\|_{2}^{2}$ has hessian $A-\lambda_{\min}(\mathbf{A})\geq 0$, so it is convex, which yields that $f_{\mathbf{A}}$ is $-\lambda_{\min}(\mathbf{A})$-weakly convex. ∎
@@
$-\lambda_{\max}\left(\tfrac{1}{2n}I_{n}\otimes\mathbf{D}_{\mathbf{X}}\right)\geq-\tfrac{1}{2n}\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}$
@@ C.2 Stability Analysis
[…]
@@ Appendix D Additional Empirical Results
In this section, we report additional empirical results revolving around the regularization of encoder. First in  D.1, we conduct the stability analysis when regularizing the encoder. Then, we report further results of encoder and decoder regularization on DSprites. Lastly, we take a first exploratory step towards decoder-free disentangled representation learning in D.3.
@@ D.1 Encoder Stability Analysis
[…]
@@ D.2 Encoder Analysis on DSprites
DSprites Higgins et al. (2017)
Base $27.6$$\pm 13.4$ $36.0$$\pm 5.3$ 38.7 ±$\pm 15.7$ $48.1$$\pm 10.8$
+ Enc-DST 32.8 ±$\pm 15.0$ $36.5$$\pm 5.9$ $33.9$$\pm 15.9$ $48.9$$\pm 11.1$
+ Enc-GMG $27.5$$\pm 14.3$ 37.4 ±$\pm 5.8$ $31.0$$\pm 14.3$ $45.9$$\pm 10.9$
@@
+ Dec-GMG 39.5 ±$\pm 15.2$ 42.2 ±$\pm 3.6$ 46.7 ±$\pm 2.0$ 50.1 ±$\pm 8.5$
D.3 Towards Decoder-free Disentanglement
@@
[…]
Recently, works such as Burns et al. (2021), von Kügelgen et al. (2021), Eastwood et al. (2023), Matthes et al. (2023), Aitchison and Ganev (2024) have shown the possibility of disentanglement through self-supervised, contrastive learning objectives in an effort to align with the scalability of encoder-only representation learning (Chen et al., 2020b, Zbontar et al., 2021, Bardes et al., 2022, Garrido et al., 2023). However, these encoder-only approaches still require weak supervision or access to multiple views of an image to learn meaningful representations of the data samples.

### v3

@@ B.2 On rescaling the costs matrices in the entropic GW solver
[…]
@@ B.3 Positivity of the Entropic GMG estimator
[…]
For any coupling $\mathbf{P}\in U_{n}$, since $\varepsilon H(\mathbf{P})=-\varepsilon\sum_{i,j=1}^{n}\mathbf{P}_{ij}\log(\mathbf{P}_{ij})<0$, one has:
@@
B.5 Proofs of Thm. 3.4
See 3.4
Before proving Thm. 3.4, we first demonstrate some technical results that will be useful later.
@@
Let ${p}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$ and ${q}_{n}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{y}_{i}}$ two uniform, empirical measures, supported on the same number of points. We denote by $P_{n}=\{\mathbf{P}\in\mathbb{R}^{n\times n},\exists\sigma\in\mathcal{S}_{n},\mathbf{P}_{ij}:=\delta_{j,\sigma(i)}\}$ the set set of permutation matrices. Assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$) are conditionally positive kernels (see A.1). Then, the GM and GW formulations coincide, in the sense that we can restrict the GW problem to permutations, namely
@@
As we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$) are conditionally positive kernels, from (Séjourné et al., 2023, Theorem. 3), $\mathbf{P}^{\star}$ also solves:
@@
[…]
@@
We use the fact that a twice continuously differentiable function is convex i.f.f. its hessian is positive semi-definite (Boyd and Vandenberghe, 2004, §(3.1.4)). Therefore, $f_{\mathbf{A}}$ is convex i.f.f. $\nabla^{2}f_{\mathbf{A}}=\mathbf{A}\geq 0$. If $\lambda_{\min}(\mathbf{A})\geq 0$, then $\mathbf{A}\geq 0$ so $f_{\mathbf{A}}$ is convex, i.e. 0-weakly convex. Otherwise, $f_{\mathbf{A}}-\tfrac{1}{2}\lambda_{\min}(\mathbf{A})\|\cdot\|_{2}^{2}$ has hessian $A-\lambda_{\min}(\mathbf{A})\geq 0$, so it is convex, which yields that $f_{\mathbf{A}}$ is $\lambda_{\min}(\mathbf{A})-$weakly convex. ∎
@@
$\lambda_{\max}\left(\tfrac{1}{2n}I_{n}\otimes\mathbf{D}_{\mathbf{X}}\right)\geq-\tfrac{1}{2n}\max_{i=1,\dots,n}\|\mathbf{x}_{i}\|_{2}^{2}$
@@ C.2 Stability Analysis
For the gradient stability analysis experiment, we repeat the following experiment for each of the four image datasets $\mathcal{D}$ that we consider. We first considered a fixed neural map $T_{\theta}$, which we choose to be a randomly initialized neural network, consisting of encoder and decoder, before any training. For $k=1,\dots,5$, we sample a batch $\mathbf{x}_{1}^{k},\dots,\mathbf{x}_{n}^{k}\sim\mathcal{D}$ and let $r_{n}^{k}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}^{k}}$. We report the pariwise cosine similarity between the gradients of the DST and the GMG. Formally, we compute $\textrm{cos-sim}(\nabla_{\theta}\mathrm{DST}_{r_{n}^{k}}(T_{\theta}),\nabla_{\theta}\mathrm{DST}_{r_{n}^{l}}(T_{\theta}))$, and $\textrm{cos-sim}(\nabla_{\theta}\mathrm{GMG}_{r_{n}^{k}}(T_{\theta}),\nabla_{\theta}\mathrm{GMG}_{r_{n}^{l}}(T_{\theta}))$, for $k,l=1,\dots,5$.
@@ Appendix D Additional Empirical Results
In this section, we report additional empirical results revolving around the regularization of encoder. First in  D.1, we conduct the stability analysis when regularizing the encoder. Then, we report further results of encoder and decoder regularization on DSprites. Lastly, we take a first exploratory step towards decoder-free disentangled representation learning in D.6.
@@ D.1 Encoder Stability Analysis
We repeat the following experiment for each of the four image datasets $\mathcal{D}$ that we consider. We first considered a fixed neural map $T_{\theta}$. For $i=1,\dots,5$, we sample a batch $\mathbf{x}_{1}^{i},\dots,\mathbf{x}_{n}^{i}\sim\mathcal{D}$ and let $r_{n}^{i}=\tfrac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{n}^{i}}$. We report the pariwise cosine similarity between the gradients of the DST. Namely, for $i,j=1,\dots,5$, we compute $\textrm{cos-sim}(\nabla_{\theta}\mathrm{DST}_{r_{n}^{i}}(T_{\theta}),\nabla_{\theta}\mathrm{DST}_{r_{n}^{j}}(T_{\theta}))$, and the GMG, $\textrm{cos-sim}(\nabla_{\theta}\mathrm{GMG}_{r_{n}^{i}}(T_{\theta}),\nabla_{\theta}\mathrm{GMG}_{r_{n}^{j}}(T_{\theta}))$.
@@ D.2 Encoder Analysis on DSprites
DSprites (Higgins et al., 2017)
Base $27.6$ $\pm 13.4$ $36.0$ $\pm 5.3$ 38.7 $\pm 15.7$ $48.1$ $\pm 10.8$
+ Enc-DST 32.8 $\pm 15.0$ $36.5$ $\pm 5.9$ $33.9$ $\pm 15.9$ $48.9$ $\pm 11.1$
+ Enc-GMG $27.5$ $\pm 14.3$ 37.4 $\pm 5.8$ $31.0$ $\pm 14.3$ $45.9$ $\pm 10.9$
@@
+ Dec-GMG 39.5 $\pm 15.2$ 42.2 $\pm 3.6$ 46.7 $\pm 2.0$ 50.1 $\pm 8.5$
D.3 Ablation on the Entropic Regularization Strength
Effect of $\varepsilon$ on the entropic GW solver.
In this section, we remind some insights about the entropic GW solver (Peyré et al., 2016) introduced in § 2.2 and used to compute the GMG in Alg. 1. We provide its algorithmic details in Alg. 2. This solver naturally provides a trade-off between (i) the approximation of the true GW distance $\mathrm{GW}(p,q)$ and (i) the optimization speed (i.e., convergence rate). This trade-off is controlled by the entropic regularization strength $\varepsilon$.
  • (i)
    Optimization. The solver employs a mirror descent scheme that iteratively linearizes the entropic GW problem (EGWP) and applies the Sinkhorn algorithm. Each mirror descent step corresponds to a projection with respect to the KL divergence, which can be efficiently performed using the Sinkhorn algorithm. The solver is initialized with $\mathbf{P}_{0}=\tfrac{1}{n^{2}}\mathbf{1}\mathbf{1}^{\top}$ and iterates as $\mathbf{P}_{t+1}\leftarrow\textsc{Sinkhorn}\left(-\mathbf{C}_{\mathcal{X}}\mathbf{P}_{t}\mathbf{C}_{\mathcal{Y}},\varepsilon\right)$. Since a larger $\varepsilon$ in Sinkhorn leads to faster convergence (Cuturi, 2013, Altschuler et al., 2018), this subsequently accelerates each mirror descent step. In other words, it reduces the number of iterations within each Sinkhorn call, i.e., the number of inner iterations of the solver. Furthermore, Rioux et al. (2023) recently demonstrated that increasing $\varepsilon$ enhances the convexity of the entropic GW problem, thereby improving the convergence rate of the mirror descent scheme. This provides theoretical justification for using a larger $\varepsilon$ to reduce the number of mirror descent steps. In other words, it reduces the number of calls to Sinkhorn, i.e., the number of outer iterations of the solver.
    We have empirically validated this behavior on DSprites with BetaVAE + GMG. We plot the mean amount of inner and outer iterations for the first 3 epochs over 5 seeds for six different values of $\varepsilon_{0}$, which we provide in Figure 5. We can observe the expected scaling of increased inner and outer iterations with decreased $\varepsilon_{0}$.
  • (ii)
    Approximation. Zhang et al. (2023) show that when using inner products and squared Euclidean distances as costs, and for $\varepsilon\in(0,1]$, the approximation error scales as:
    $|\mathrm{GW}_{\varepsilon}(p,q)-\mathrm{GW}(p,q)|\lesssim\varepsilon\log(1/\varepsilon),$
    where the constants in $\lesssim$ depend on $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$, that is, the dimensions of the support of $p$ and $q$, as well as their fourth-order moments.
Given this trade-off, the practical goal is to select an $\varepsilon$ value that is sufficiently large to ensure fast convergence while avoiding any degradation in performance. Across all our experiments, we found that $\varepsilon=0.1$ struck the right balance. We validate this observation in the next paragraph.
Algorithm 2 Entropic Gromov-Wasserstein solver (Peyré et al., 2016), (Scetbon et al., 2022, Alg. 2).
1:Require: samples $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\sim p$; $\mathbf{y}_{1},\dots,\mathbf{y}_{n}\sim q$; cost functions $c_{\mathcal{X}},c_{\mathcal{Y}}$; entropic regularization scale $\varepsilon_{0}$ (default =0.1=0.1), statistic operator on cost matrix stat (default $\texttt{mean}$).
[…]
6:while converged do
7:  $\mathbf{C}_{t+1}\leftarrow-\mathbf{C}_{\mathcal{X}}\mathbf{P}_{t}\mathbf{C}_{\mathcal{Y}}$ $\triangleright$ $n^{3}$ or $n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}})$
8:  $\mathbf{P}_{t+1}\leftarrow\textsc{Sinkhorn}(\mathbf{C}_{t+1},\varepsilon)$ $\triangleright$ $\mathcal{O}(n^{2})$
9:end while
10:Compute $\mathrm{GW}_{\varepsilon}(p_{n},q_{n})$ from $\mathbf{P}_{t}$ using Eq. (EGWP) $\triangleright$ $n^{3}$ or $n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}})$
11:return $\mathrm{GW}_{\varepsilon}(p_{n},q_{n})$
Effect of $\varepsilon$ in disentanglement.
We investigate the effect of the entropic regularization strength $\varepsilon$ used to compute GMG on the disentanglement performances. The results are presented in Fig. 5 and show that performance is robust to the choice of entropic regularization scale $\varepsilon_{0}$. We observed this both with respect to a setting, where we see major improvements from the GMG (DSprites with BetaVAE) as well as one where we only observe minor improvements (DSprites with BetaTCVAE). This validates our choice to use a single reasonable value for all our experiments, namely $\varepsilon_{0}=0.1$. This robustness with respect to the entropic strength was also observed in a recent work (Piran et al., 2024) proposing a similar gap regularization, based on the entropic multi-marginal OT problem. See (Piran et al., 2024, Fig. 3) for experiments highlighting this robustness. Our intuition is as follows: when using Sinkhorn (in the case of M3G) or GW/quadratic OT (in our work) to compute a training loss, $\varepsilon$ acts as a sharpness parameter, emphasizing certain pairs of points more strongly. While the loss value changes with $\varepsilon$, the optimization of network variables on top of this loss appears to be largely unaffected by the sharpness introduced by $\varepsilon$. However, when the output of Sinkhorn or GW is used directly for predictions or learning flows, for example, in Monge maps (Pooladian and Niles-Weed, 2021, Kassraie et al., 2024) or Gromov-Monge maps (Klein et al., 2024), the entropic regularization strength $\varepsilon$ has a much stronger influence and requires careful tuning.
Figure 5: Analysis of the effect of the entropic regularization scale $\varepsilon_{0}$ on the disentanglement performances when learning with the GMG. We use the SmallNORB dataset and consider the $\beta$-VAE (in blue) and $\beta$-TCVAE (in orange) settings. The GMG is applied to the decoder $d_{\theta}$ with cost functions $c_{\mathcal{X}}=c_{\mathcal{Y}}=\mathrm{cos}\text{-}\mathrm{sim}(\cdot)$, corresponding to the setting that provides better disentanglement. We compute the GMG using Algorithm 1 with stat=mean. We investigate the effect of the entropic regularization scale $\varepsilon_{0}$ by testing five other values of $\varepsilon_{0}$ besides the one used in all other experiments in the paper (namely, $\varepsilon_{0}=0.1$). The values tested are $\varepsilon_{0}\in\{0.05,0.2,0.3,0.5,1\}$. We also include the baseline result without using the GMG (dashed line) as a comparison. Additionally, we provide a runtime analysis with respect to both the inner and outer iteration of the GW-solver.
D.4 Scaling to Higher Image Resolutions
As detailed in § 3.2, the computation of the GMG scales linearly with the data dimension, enabling our method to handle high-dimensional settings effectively. To demonstrate this scalability, we benchmark our approach on the Shapes3D dataset upscaled to $128\times 128$ resolution. To accommodate the increased image resolution, we extend the Decoder by adding one additional layer. Using our best-performing configuration from the $64\times 64$ experiments—Cos costs and the GMG applied to the decoder $d_{\theta}$—we conduct a focused hyperparameter search as described in Table 5. The results, summarized in Table 6, compare the four base models Beta(TC)VAE (+ HFS), both with and without the GMG. Our findings indicate that the proposed setup scales to higher resolutions while preserving the performance improvements achieved by the GMG. Finally, we visually validate these results by plotting the latent traversals for the best-performing configuration, as shown in Figure 6.
Table 5: Hyperparameter grid search for Shapes3D $128\times 128$.
Method Parameter Values
$\beta$-(TC)VAE $\beta$ [10, 16]
+ GMG $\lambda$ [0.1, 1, 10]
$\beta$-(TC)VAE $\beta$ [2, 4]
+ GMG $\lambda$ [0.1, 1, 10]
+ HFS $\gamma$ [1, 10]
Table 6: Impact of the GMG, applied with $\mathbf{Cos}$ as the cost function on the decoder $d_{\theta}$, on disentanglement performance using upscaled $128\times 128$ Shapes3D images. Performance is evaluated using DCI-D, with the best result highlighted for each method.
Shapes3D ($128\times 128$) (Burgess and Kim, 2018)
Base $54.3$ $\pm 19.3$ 74.6{74.6} $\pm 17.3$ $87.2$ $\pm 2.7$ $88.6$ $\pm 11.4$
+ Dec-GMG 63.9 $\pm 8.2$ 82.0 $\pm 12.8$ 90.4 $\pm 3.7$ 92.2 $\pm 8.2$
D.5 Latent Traversal Visualization
Figure 6: Latent traversal visualization for Shapes3D $128\times 128$ with our best performing setup, BetaTCVAE + HFS + GMG. We select the best performing result out of 5 seeds achieving a DCI-D of $99.4$. We plot four different latent dimensions while traversing them from 1.0-1.0 to $1.0$. As visualized the model has clearly learned to separate wall hue, object hue, scale, and floor hue into different latent dimensions.
Figure 7: Latent traversal visualization for Shapes3D $64\times 64$ with our best performing setup, BetaTCVAE + HFS + GMG. We select the best performing result out of 5 seeds achieving a DCI-D of $100.0$. We plot four different latent dimensions while traversing them from 1.0-1.0 to $1.0$. As visualized the model has clearly learned to separate wall hue, object hue, scale, and floor hue into different latent dimensions.
D.6 Towards Decoder-free Disentanglement
@@
Decoder-free $\beta$-VAE $\beta$-TCVAE Base $0.0$ $\pm 0.0$ $0.0$ $\pm 0.0$ $\mathbf{L2^{2}}:\|\cdot-\cdot\|_{2}^{2}$ + DST $38.2$ $\pm 0.8$ $42.7$ $\pm 1.6$ + GMG $13.9$ $\pm 0.4$ $20.5$ $\pm 0.5$ $\mathbf{ScL2^{2}}:\alpha\|\cdot-\cdot\|_{2}^{2}$, $\alpha>0$ learnable + DST 45.6 $\pm 1.2$ 53.5 $\pm 1.0$ + GMG $15.2$ $\pm 0.3$ $25.2$ $\pm 0.6$ $\mathbf{Cos}:\textrm{cos-sim}(\cdot,\cdot)$ + DST $37.0$ $\pm 0.4$ $46.1$ $\pm 1.5$ + GMG $37.0$ $\pm 0.9$ $38.8$ $\pm 1.1$
Recently, works such as (Burns et al., 2021, von Kügelgen et al., 2021, Eastwood et al., 2023, Matthes et al., 2023, Aitchison and Ganev, 2024) have shown the possibility of disentanglement through self-supervised, contrastive learning objectives in an effort to align with the scalability of encoder-only representation learning (Chen et al., 2020b, Zbontar et al., 2021, Bardes et al., 2022, Garrido et al., 2023). However, these encoder-only approaches still require weak supervision or access to multiple views of an image to learn meaningful representations of the data samples.